OpenStudy (♥ashley♥):

Q is the midpoint of GH¯¯¯¯¯¯ , GQ=2x+3, and GH=5x−5 . What is the length of GQ¯¯¯¯¯ ? Enter your answer in the box. I know the answer is 25 but why and how?

563blackghost (563blackghost):

It'd be best to draw it out to understand.... 

563blackghost (563blackghost):

We see that we know of the equation to equal the entire length and we know GQ which is half of the length is 2x + 3 so in order to find GQ we would need to find the equation of GH....So we would subtract 5x - 5 and 2x + 3.... \(\Huge{(5x-5)-(2x+3)=GQ}\) So what would that simplify to?

OpenStudy (♥ashley♥):

x=8/3

OpenStudy (alivejeremy):

O_O

563blackghost (563blackghost):

So so sorry my internet is bad...I meant we need to find QH in which the equation above is to find QH

563blackghost (563blackghost):

So... \(\huge{(5x-5)-(2x+3)=QH}\) We first combine like terms to solve... \(\huge{3x-8=QH}\) So we know for a fact that QH and GQ have equal lengths in which we can use to find x.... \(\huge{3x-8=2x+3}\) What would this simplify to?

OpenStudy (needshelpdfdf):

ashley

OpenStudy (♥ashley♥):

x=11

OpenStudy (♥ashley♥):

Then it both sides equals to 25

563blackghost (563blackghost):

\(\huge{\checkmark}\)

563blackghost (563blackghost):

Nicely done!

OpenStudy (♥ashley♥):

Thank you for your help!

563blackghost (563blackghost):

np :) sorry for long and slow replies my internet is bad

OpenStudy (♥ashley♥):

It's okay :)